OpenStudy (anonymous):

write an equation for the line perpendicular to the given line that contains p. P(1,1); y=4x+3

OpenStudy (anonymous):

slope of \[\huge y=\color{blue}mx+b\] is \[\huge \color{blue}m\] what is the slops of \[y=4x+3\]?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

would it be y-1=4(x-1)

OpenStudy (anonymous):

ok good now the "perpendicular" line has a slope that is the negative recrocal \[\huge -\frac{1}{\color{blue}m}\]

OpenStudy (anonymous):

"reciprocal"

OpenStudy (anonymous):

what is the slope of the perpendicular line?

OpenStudy (anonymous):

i mean the slope would be -1/4

OpenStudy (anonymous):

right, so \[y-1=-\frac{1}{4}(x-1)\] is a start

OpenStudy (anonymous):

is that all i need?

OpenStudy (mathmale):

What do you think? The above equation is one form of the desired line. You could solve this equation for y and end up with the slope-intercept form.

OpenStudy (anonymous):

the instructions don't say to put in y intercept it just says to leave it! thank you very much !

OpenStudy (anonymous):

satellite knows the following: The slope (m) x the perpenicular slope(1/m) = -1 its a basic rule in maths.

OpenStudy (mathmale):

If you have \[y-1=-\frac{1}{4}(x-1)\] we say this equation is in point-slope form. Were you to solve for y, you'd get the equivalent equation in slope-intercept form. (You don't "put in y intercept;" rather you IDENTIFY the y-intercept as the constant term on the right: y = mx + b (b is the y-intercept).